# TinySAM: Pushing the Envelope for Efficient Segment Anything Model

### 21 Dec 2023 v1

@@ 5 Conclusion
In this paper, we propose a framework to push the envelope for segment anything task and obtain a highly efficient model named TinySAM. We first propose a full-stage knowledge distillation method with online hard prompt sampling strategy to distill a lightweight student model. We also adapt the post-training quantization to the promptable segmentation task and further reducing the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and push the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterparts methods. We hope the proposed TinySAM brings beneficial perspective for designing a highly efficient segment anything model.

### 09 Mar 2024 v2

@@ 5 Conclusion
In this paper, we propose a framework to push the envelope for segment anything task and obtain a highly efficient model named TinySAM. We first propose a full-stage knowledge distillation method with hard mask weighting and hard prompt sampling strategy to distill a lightweight student model. We also adapt the post-training quantization to the promptable segmentation task and further reducing the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and push the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterparts methods. We hope the proposed TinySAM brings beneficial perspective for designing a highly efficient segment anything model.